# Preference-based Centrality and Ranking in General Metric Spaces

### v1

@@ 3 Theoretical Properties
This section studies the population properties of the calibrated centrality score $\theta^{\star}$ defined through the M-projection in (4). Our analysis proceeds in three steps. First, we establish that the population projection problem is well-posed and admits a unique solution. Second, we show that the resulting score is intrinsically linked to the preference-based centrality functional $r(\cdot)$ through a monotone transformation, independently of correct model specification. Finally, we demonstrate that the induced population score satisfies the canonical axioms of statistical depth and recovers familiar center–outward orderings in symmetric benchmark settings.
@@ 3.1 Existence and uniqueness of the centered minimizer
We begin by showing that the population M-projection defining $\theta^{\star}$ is well-defined. Since the optimization is carried out over a space of functions, mild regularity conditions are required to ensure existence and identifiability of the solution.
Assumption 1 (Centered parameter class).
@@
Assumption 1 is standard in functional M-estimation. Convexity rules out multiple local optima, while compactness guarantees existence of a minimizer in the infinite-dimensional function space.
@@
Under Assumption 1, the risk functional $\mathcal{L}_{F}$ defined in (3) has a unique minimizer over $\Theta_{0}(F)$.
Theorem 1 ensures that the M-projection problem admits a unique population solution $\theta^{\star}\in\Theta_{0}(F)$. This result identifies a single, well-defined calibrated score function that serves as the population target of the proposed methodology.
@@
If the working BTL model happens to be correctly specified, so that $p_{F}(X,Y)=q_{\theta^{\star}}(X,Y)$ almost surely, then $\theta^{\star}$ coincides with the true underlying score function. In this case, the M-projection reduces to the usual likelihood-based identification, with the centering constraint selecting a unique representative among shift-equivalent solutions.
3.2 A monotone link between calibrated scores and preference centrality
The existence and uniqueness result above ensures that the calibrated score $\theta^{\star}$ is well-defined. We now show that this score is intrinsically linked to the preference-based centrality functional $r(\cdot)$ introduced in Section 1. Importantly, this link holds without assuming correct specification of the BTL working model.
Theorem 2 (Monotone calibration link between $\theta^{\star}$ and $r$).
Suppose Assumption 1 holds and that $\theta^{\star}$ lies in the relative interior of $\Theta_{0}(F)$. Define
@@ 3.2 A monotone link between calibrated scores and preference centrality
Then $H:(-\infty,\infty)\to(0,1)$ is strictly increasing and satisfies
$H\big(\theta^{\star}(y)\big)=r(y)\qquad\text{for $F$-a.e.\ }y\in\mathcal{X}.$ (11)
In particular, $\theta^{\star}(\cdot)$ is a strictly monotone transformation of $r(\cdot)$, and
$\theta^{\star}(y)=H^{-1}\big(r(y)\big)\qquad\text{for $F$-a.e.\ }y\in\mathcal{X},$
@@
Theorem 2 establishes that the calibrated score $\theta^{\star}(\cdot)$ is a strictly monotone reparameterization of the preference centrality $r(\cdot)$. As a consequence, $\theta^{\star}$ and $r$ induce identical population orderings up to $F$-null sets, and the M-projection preserves the intrinsic preference structure while mapping it onto a coherent one-dimensional scale.
The relative-interior condition is a mild technical assumption ensuring that first-order optimality conditions characterize the solution of the constrained optimization problem. It rules out boundary solutions at which only restricted perturbations are admissible.
@@ 3.3 Depth properties of the M-projection scores
While $\theta^{\star}(\cdot)$ provides a calibrated ranking, it is natural to ask whether the induced strength score $s^{\star}(\cdot)=\exp\{\theta^{\star}(\cdot)\}$ admits a genuine center–outward interpretation. We answer this question by showing that, under mild regularity conditions, $s^{\star}$ satisfies the four canonical axioms of statistical depth (Zuo and Serfling, 2000).
Assumption 2 (Center–outward regularity of preference centrality).
Assume that the preference-based centrality functional $r(\cdot)$ satisfies the following population-level regularity conditions.
  1. (A1)
    Affine equivariance. For any invertible matrix $A$ and vector $b$, define the affine transformation $T_{A,b}(x)=Ax+b$ and let $G=T_{A,b}\#F$ denote the pushforward distribution. The induced preferences are equivariant in the sense that
    $p_{G}\bigl(T_{A,b}x,T_{A,b}y\bigr)=p_{F}(x,y)\quad\text{for all }x,y\in\mathcal{X},$
    and the centered parameter class transforms accordingly as
@@
  2. (A2)
    Existence of a preference center and center–outward monotonicity. There exists at least one maximizer $\mu\in\arg\max_{y\in\mathcal{X}}r(y)$ such that for any $x\in\mathcal{X}$ and any $\alpha\in[0,1]$, the interpolated point $x_{\alpha}=(1-\alpha)\mu+\alpha x$ satisfies
    $r(x_{\alpha})\geq r(x).$
    When $\arg\max r$ is not a singleton, we fix any maximizer $\mu$ satisfying this property.
  3. (A3)
    Vanishing of centrality at infinity. Let $Z\sim F$. For any sequence $(y_{m})$ with $\delta(y_{m},\mu)\to\infty$ and any fixed radius $M>0$,
    $\mathbb{P}\{\delta(y_{m},Z)\leq M\}\to 0,$
    so that points arbitrarily far from the center have asymptotically negligible preference centrality.
Assumption 2 imposes population-level center–outward regularity on the preference centrality functional $r(\cdot)$. Collectively, conditions (A1)–(A3) mirror the canonical depth axioms of Zuo and Serfling (2000): affine equivariance, maximality at a center, monotonicity toward the center, and vanishing at infinity. Via the monotone link in Theorem 2, these properties transfer directly from $r(\cdot)$ to the CORE score $s^{\star}(\cdot)$.
Theorem 3 (Depth properties).
Suppose Assumptions 1 and 2 hold and that $\theta^{\star}$ lies in the relative interior of $\Theta_{0}(F)$. Then $s^{\star}(x)=\exp\{\theta^{\star}(x)\}$ satisfies the four canonical depth axioms:
@@
    Affine invariance. For $G=T_{A,b}\#F$,
    $s_{G}^{\star}\big(T_{A,b}x\big)=s^{\star}(x).$
@@
    Maximality at the center. If $\mu$ is as in Assumption 2(A2), then $s^{\star}(\mu)\geq s^{\star}(x)$ for all $x$.
@@
    Monotonicity toward the deepest point. For $x_{\alpha}=(1-\alpha)\mu+\alpha x$ with $\alpha\in[0,1]$,
    $s^{\star}(x_{\alpha})\geq s^{\star}(x).$
@@
    Vanishing at infinity. If $\delta(y_{m},\mu)\to\infty$, then $s^{\star}(y_{m})\to 0$.
Theorem 3 provides an axiomatic characterization of the CORE population score $s^{\star}$ as a statistical depth function. While depth axioms describe qualitative behavior, it is also instructive to examine the explicit form of $\theta^{\star}$ in symmetric benchmark models.
@@ 3.3.1 One-dimensional symmetric distribution
We begin with the one-dimensional Euclidean setting $\mathcal{X}=\mathbb{R}$. In this case, symmetry about a center $\mu$ suggests that any reasonable population centrality should assign equal importance to $\mu+t$ and $\mu-t$ and depend on $x$ only through its deviation from the center. This intuition underlies classical univariate depth notions and motivates the reflection-symmetry assumption below.
@@
Assume $\mathcal{X}=\mathbb{R}$ and $\delta(x,y)=|x-y|$. Assume that for $X\sim F$,
$X-\mu\ \overset{d}{=}\ \mu-X,$
so that $F$ is symmetric about $\mu$.
@@
Suppose the conditions of Theorem 2 are satisfied and Assumption 3 holds. Then there exists a function $\varphi\colon[0,\infty)\to\mathbb{R}$ such that
@@
If in addition Assumption 2(A2) holds, then $\varphi$ can be chosen non-increasing. In particular, if $|x-\mu|<|y-\mu|$, then $\theta^{\star}(x)\geq\theta^{\star}(y)$.
Proposition 1 shows that, under reflection symmetry, the calibrated population score depends only on the distance to the center and decreases monotonically as one moves away from $\mu$. Consequently, $\theta^{\star}$ is maximized at the symmetry center $\mu$ and induces the canonical one-dimensional center–outward ordering. Importantly, this structural characterization holds regardless of whether the BTL working model correctly specifies the underlying preference probabilities, demonstrating that the proposed calibration preserves the intrinsic geometry implied by distributional symmetry.
@@ 3.3.2 Multivariate elliptically symmetric distribution
We next consider the multivariate Euclidean setting $\mathcal{X}=\mathbb{R}^{d}$ under elliptical symmetry. In this setting, the geometry of the distribution is characterized by a location parameter $\mu$ and a positive definite scatter matrix $\Sigma$, and centrality is naturally described in terms of distances measured relative to $\Sigma$. Classical multivariate depth constructions, most notably Mahalanobis depth, exploit this structure to induce center–outward orderings based on ellipsoidal level sets.
To make this connection precise, we equip $\mathcal{X}$ with the Mahalanobis dissimilarity associated with $\Sigma$ and formalize elliptical symmetry through an invariance condition in whitened coordinates.
@@
$\delta_{\Sigma}(x,y):=\bigl\{(x-y)^{\top}\Sigma^{-1}(x-y)\bigr\}^{1/2},\qquad\Sigma\in\mathbb{R}^{d\times d}\ \text{positive definite}.$
Assume that for $X\sim F$ and every orthogonal matrix $O$,
@@
which characterizes elliptical symmetry of $F$ about $\mu$ with scatter matrix $\Sigma$.
@@
Suppose the conditions of Theorem 2 are satisfied and Assumption 4 holds. Then there exists a function $\varphi\colon[0,\infty)\to\mathbb{R}$ such that
@@
If in addition Assumption 2(A2) holds, then $\varphi$ can be chosen non-increasing. In particular, if $\delta_{\Sigma}(x,\mu)<\delta_{\Sigma}(y,\mu)$, then $\theta^{\star}(x)\geq\theta^{\star}(y)$.
Proposition 2 shows that, under elliptical symmetry, the CORE population score induces the same center–outward ordering as classical Mahalanobis depth. Specifically, the induced centrality score is constant on ellipsoids centered at $\mu$ with shape matrix $\Sigma$ and decreases monotonically with the Mahalanobis radius $\delta_{\Sigma}(x,\mu)$. Consequently, the score is maximized at the location parameter $\mu$ and recovers the canonical multivariate center–outward ordering characteristic of Mahalanobis depth. This confirms that our framework correctly adapts to the underlying covariance structure, effectively accounting for the geometry encoded by $\Sigma$ to recover the intrinsic notion of multivariate centrality.

### v2

@@ 3 Theoretical Properties
This section establishes population and sample-level guarantees for the calibrated centrality score $\theta^{\star}$ defined in (4). Our results clarify (i) when the population M-projection is well-posed and uniquely defined, (ii) how the resulting score relates to the intrinsic preference centrality $r_{F}(\cdot)$ without requiring correct specification of the BTL working model, and (iii) what structural and statistical properties the calibrated score inherits, including depth axioms, symmetry-induced orderings, and consistency of the sample estimator.
@@ 3.1 Existence and uniqueness of the centered minimizer
We begin by establishing that the population M-projection defining $\theta^{\star}$ is well-defined. Because the optimization is carried out over a function class, we impose a mild compactness and convexity condition to guarantee existence and rule out non-identifiability beyond the intrinsic shift invariance handled by centering.
Assumption 1.
@@
Assumption 1 is standard in functional M-estimation. Compactness ensures existence of a minimizer in the infinite-dimensional parameter space, and convexity supports uniqueness once the objective is shown to be strictly convex on $\Theta_{0}(F)$.
@@
Under Assumption 1, the risk functional $\mathcal{L}_{F}$ defined in (3) is strictly convex on $\Theta_{0}(F)$ and therefore admits a unique minimizer over $\Theta_{0}(F)$ up to $F$-a.e. equality, where $F$-a.e. means that the stated property holds outside a measurable set $\mathcal{Q}\subset\mathcal{X}$ with $\mathbb{P}_{F}(\mathcal{Q})=0$.
Theorem 1 identifies a single, well-defined calibrated score function $\theta^{\star}\in\Theta_{0}(F)$ as the population target of our methodology.
@@
If the working BTL form is correctly specified in the sense that $p_{F}(X,Y)=q_{\theta^{\star}}(X,Y)$ almost surely for some centered score $\theta^{\star}\in\Theta_{0}(F)$, then the M-projection recovers that underlying score function (up to the centering convention).
3.2 The monotone calibration link
Having established that $\theta^{\star}$ is uniquely defined, we next show that it provides an order-preserving reparameterization of the preference centrality $r_{F}(\cdot)$ introduced in Section 1. This link is intrinsic to the projection construction and does not rely on correctness of the BTL working model.
Assumption 2.
The calibrated score $\theta^{\star}$ lies in the relative interior of $\Theta_{0}(F)$.
Theorem 2 (Monotone calibration link between $\theta^{\star}$ and $r_{F}$).
Define
@@ 3.2 The monotone calibration link
Under Assumptions 1 and 2, $H:(-\infty,\infty)\to(0,1)$ is strictly increasing and satisfies
$H\big(\theta^{\star}(x)\big)=r_{F}(x)\qquad\text{for $F$-a.e.\ }x\in\mathcal{X}.$ (11)
Consequently, $\theta^{\star}(\cdot)$ is a strictly monotone transformation of $r_{F}(\cdot)$, and
$\theta^{\star}(x)=H^{-1}\big(r_{F}(x)\big)\qquad\text{for $F$-a.e.\ }x\in\mathcal{X},$
@@
Theorem 2 implies that $\theta^{\star}$ and $r_{F}$ induce the same population ordering up to $F$-null sets: the M-projection preserves the intrinsic preference-based ranking while providing a coherent one-dimensional score scale. Assumption 2 is a technical condition ensuring that the constrained minimizer is characterized by first-order optimality conditions (i.e., excluding boundary solutions where only restricted perturbations are admissible).
The M-projection defines $\theta^{\star}$ as the unique centered minimizer of a global logistic cross-entropy risk. Equivalently, it minimizes the integrated Kullback–Leibler divergence between the intrinsic preference kernel $p_{F}(x,y)$ and its BTL-type approximation $q_{\theta}(x,y)=\sigma\{\theta(x)-\theta(y)\}$. Thus, $\theta^{\star}$ is the closest logit-consistent representation of $p_{F}$ within the BTL family.
Although $\theta^{\star}$ and $r_{F}$ induce identical population orderings, $\theta^{\star}$ additionally provides a coherent log-odds scale satisfying
$q_{\theta^{\star}}(x,y)=\sigma\{\theta^{\star}(x)-\theta^{\star}(y)\}.$
This probabilistic representation turns preference aggregation into a convex M-projection problem with a unique target, which facilitates estimation and finite-sample analysis.
@@ 3.3 Depth properties of the M-projection scores
The previous subsection established that the calibrated score $\theta^{\star}$ is uniquely defined and is a strictly monotone reparameterization of the intrinsic preference centrality $r_{F}(\cdot)$. We now investigate whether the logistic M-projection preserves the center–outward geometry encoded in the underlying preference relation.
Statistical depth provides an axiomatic framework for formalizing center–outward structure. We adopt the classical affine depth axioms of Zuo and Serfling (2000). Since these axioms are formulated through affine transformations and linear interpolation, we restrict in this subsection to the Euclidean setting $\mathcal{X}=\mathbb{R}^{d}$ and treat $F$ as a distribution on $\mathbb{R}^{d}$. The dissimilarity $\delta$ that defines the preference kernel $p_{F}(x,y)$ is allowed to be any measurable dissimilarity on $\mathbb{R}^{d}$; it need not coincide with the Euclidean norm.
Rather than assuming that $r_{F}(\cdot)$ is itself a depth function, we impose structural regularity directly on the preference geometry. The following conditions encode center–outward behavior at the level of the preference kernel.
Assumption 3 (Center–outward regularity of the preference geometry).
The preference kernel $p_{F}(x,y)$ satisfies:
  1. (a)
    Affine equivariance. For every invertible affine map $T_{A,b}(x)=Ax+b$ and $G=T_{A,b}\#F$,
    $p_{G}(T_{A,b}x,T_{A,b}y)=p_{F}(x,y)\quad\text{for all }x,y\in\mathbb{R}^{d},$
    and
@@
  2. (b)
    Existence of a preference center and center–outward ordering. There exists $\mu\in\mathbb{R}^{d}$ such that
    $r_{F}(\mu)=\sup_{x\in\mathbb{R}^{d}}r_{F}(x),$
    and for all $x\in\mathbb{R}^{d}$ and $\alpha\in[0,1]$,
    $r_{F}((1-\alpha)\mu+\alpha x)\geq r_{F}(x).$
  3. (c)
    Vanishing preference at $\delta$-infinity. If $\delta(y_{m},\mu)\to\infty$, then
    $r_{F}(y_{m})\to 0.$
Assumption 3 requires that the underlying preference relation exhibits affine invariance and center–outward regularity.
Assumption 4 (Euclidean regularity and continuity of $\delta$).
Assume that $\Theta_{0}(F)\subset L_{0}^{2}(F)\cap C(\mathbb{R}^{d})$, where $C(\mathbb{R}^{d})$ denotes the space of continuous real-valued functions on $\mathbb{R}^{d}$. The distribution $F$ is absolutely continuous with respect to the Lebesgue measure and has full support on $\mathbb{R}^{d}$. Moreover, for $F$-a.e. $z\in\mathbb{R}^{d}$, the map $y\mapsto\delta(z,y)$ is continuous on $\mathbb{R}^{d}$.
Assumption 4 ensures that the M-projection can be interpreted pointwise. In the $L^{2}(F)$ formulation, two scores that differ only on an $F$-null set are identified, so the value of the score at a specific location is not determined. This is problematic because several depth axioms evaluate the score at fixed spatial points, such as the center in (D2). Restricting $\Theta_{0}(F)$ to continuous functions removes this ambiguity and turns $F$-a.e. calibration identities into pointwise statements.
The continuity of $\delta$ implies that $p_{F}(x,y)$ and $r_{F}(y)$ vary continuously with their arguments. If $\delta$ were discontinuous, the induced preference structure could exhibit jump discontinuities that are incompatible with a continuous score $\theta\in\Theta_{0}(F)$. In that case, the minimizer may lie on the boundary of the parameter space, so the interior first-order calibration condition would no longer apply.
Absolute continuity and full support of $F$ ensure that the calibration relation identifies the continuous score on all of $\mathbb{R}^{d}$. Without full support, the score is not determined outside the support of $F$, and global geometric properties such as vanishing at infinity cannot be stated meaningfully.
Theorem 3 (Geometry preservation under logistic M-projection).
Suppose Assumptions 1, 2, 3 and 4 hold. Then, for all $x\in\mathbb{R}^{d}$, the calibrated score $\theta^{\star}$ satisfies:
@@
    Affine equivariance (up to centering). For $G=T_{A,b}\#F$,
    $\theta_{G}^{\star}(T_{A,b}x)=\theta_{F}^{\star}(x).$
@@
    Maximality at the center. If $\mu$ is as in Assumption 3(b), then
    $\theta^{\star}(\mu)\geq\theta^{\star}(x).$
@@
    Monotonicity toward the center. For $x_{\alpha}=(1-\alpha)\mu+\alpha x$ with $\alpha\in[0,1]$,
    $\theta^{\star}(x_{\alpha})\geq\theta^{\star}(x).$
@@
    Behavior at $\delta$-infinity. If $\delta(y_{m},\mu)\to\infty$, then
    $\theta^{\star}(y_{m})\to-\infty.$
Theorem 3 shows that the logistic M-projection preserves the intrinsic center–outward ordering encoded in the preference geometry. Although $\theta^{\star}$ is defined as the minimizer of a parametric cross-entropy risk, it retains the full affine depth structure whenever the underlying preference relation exhibits such geometry.
Since $s^{\star}(x)=\exp\{\theta^{\star}(x)\}$ is a strictly monotone transformation of $\theta^{\star}(x)$, the same ordering and center–outward properties hold for the strength score.
@@ 3.3.1 One-dimensional symmetric distribution
We now show that the regularity conditions in Assumption 3 are naturally satisfied in classical symmetric benchmark models. This clarifies that the center–outward structure assumed above is not restrictive, but instead reflects the geometry of familiar distributions.
Consider the one-dimensional Euclidean setting $\mathcal{X}=\mathbb{R}$. Under symmetry about a location parameter $\mu$, any reasonable population centrality must assign equal importance to $\mu+t$ and $\mu-t$ and depend on $x$ only through its distance from the center. This intuition underlies classical univariate depth constructions.
@@
Assume $\mathcal{X}=\mathbb{R}$ and $\delta(x,y)=|x-y|$. Suppose that $F$ is symmetric about $\mu$ such that $X\sim F$,
$X-\mu\overset{d}{=}\mu-X.$
Under this symmetry, the induced preference centrality $r_{F}(\cdot)$ is automatically maximized at $\mu$, is symmetric around $\mu$, and decreases as $|x-\mu|$ increases. Consequently, Assumption 3(b)–(c) are satisfied.
@@
Suppose Assumptions 1, 2 and 5 hold. Then there exists a function $\varphi:[0,\infty)\to\mathbb{R}$ such that
@@
If in addition Assumption 3(b) holds, $\varphi$ can be chosen non-increasing. In particular, if $|x-\mu|<|y-\mu|$, then
$\theta^{\star}(x)\geq\theta^{\star}(y).$
Proposition 1 shows that, under reflection symmetry, the calibrated population score depends only on distance to the center and decreases monotonically as one moves away from $\mu$. Thus, the M-projection preserves the intrinsic symmetry and recovers the canonical one-dimensional center–outward ordering.
@@ 3.3.2 Multivariate elliptically symmetric distribution
We next consider the multivariate Euclidean setting $\mathcal{X}=\mathbb{R}^{d}$ under elliptical symmetry. This setting generalizes reflection symmetry to higher dimensions and encompasses Gaussian and more general elliptical distributions.
Under elliptical symmetry, centrality is naturally characterized through distances measured relative to a positive definite scatter matrix $\Sigma$. Classical multivariate depth constructions, including Mahalanobis depth, induce center–outward orderings based on ellipsoidal level sets.
@@
$\delta_{\Sigma}(x,y)=\bigl\{(x-y)^{\top}\Sigma^{-1}(x-y)\bigr\}^{1/2},\qquad\Sigma\text{ positive definite}.$
Suppose that for $X\sim F$ and every orthogonal matrix $O$,
@@
so that $F$ is elliptically symmetric about $\mu$ with scatter matrix $\Sigma$.
Under this condition, the preference centrality $r_{F}(\cdot)$ is constant on ellipsoids centered at $\mu$ and decreases as the Mahalanobis radius $\delta_{\Sigma}(x,\mu)$ increases. Therefore Assumption 3(b)–(c) are again satisfied.
@@
Suppose Assumptions 1, 2 and 6 hold. Then there exists a function $\varphi:[0,\infty)\to\mathbb{R}$ such that
@@
If in addition Assumption 3(b) holds, $\varphi$ can be chosen non-increasing. In particular, if $\delta_{\Sigma}(x,\mu)<\delta_{\Sigma}(y,\mu)$, then
$\theta^{\star}(x)\geq\theta^{\star}(y).$
Proposition 2 shows that, under elliptical symmetry, the calibrated score induces the same center–outward ordering as Mahalanobis depth. Hence the logistic M-projection preserves the intrinsic geometric structure encoded by the covariance matrix $\Sigma$, rather than distorting it.
3.4 Consistency of the sample M-estimator
We now turn to statistical properties of the sample-level estimator introduced in Section 2.2. Our goal is to show that the empirical minimizer of the cross-entropy risk based on estimated pairwise preferences consistently recovers the population calibrated score $\theta^{\star}$ evaluated at the observed data points.
Importantly, the BTL form is used purely as a working projection family. No correct specification assumption is imposed; the target remains the population M-projection $\theta^{\star}$ defined in (4).
Assumption 7 (Essential boundedness of the population score).
There exists $B_{\star}<\infty$ such that
$\|\theta^{\star}\|_{L^{\infty}(F)}\leq B_{\star}.$
Assumption 7 requires boundedness only in the $F$-essential sense. It does not preclude $\theta^{\star}(x)\to$-$\infty$ as $\delta(x,\mu)\to\infty$ outside regions of non-negligible probability mass. Rather, it ensures that the population target evaluated at random draws $X\sim F$ lies in a bounded range almost surely.
This assumption is imposed solely to facilitate finite-sample analysis. It allows us to localize the empirical optimization problem to a compact parameter set, under which the logistic risk becomes uniformly strongly convex and the empirical process can be controlled via standard concentration arguments. Such localization is common in nonparametric M-estimation and does not alter the population target.
We consider two plug-in constructions of the empirical preference matrix: the leave-two-out estimator $\widehat{\mathbf{P}}^{\mathrm{lto}}$ in (7), and the reference-sample estimator $\widehat{\mathbf{P}}^{\mathrm{ref}}$ in (8) based on an independent reference sample of size $m$.
For theoretical analysis, we restrict the optimization to the bounded centered set
$\Theta_{n}(B)=\Big\{\boldsymbol{\theta}\in\mathbb{R}^{n}:\ \mathbf{1}^{\top}\boldsymbol{\theta}=0,\ \|\boldsymbol{\theta}\|_{\infty}\leq B\Big\},$
where $B>3B_{\star}$ is fixed. We define
$\hat{\boldsymbol{\theta}}^{\mathrm{lto}}\in\operatorname*{argmin}_{\boldsymbol{\theta}\in\Theta_{n}(B)}\mathcal{L}_{n}(\boldsymbol{\theta};\widehat{\mathbf{P}}^{\mathrm{lto}}),\qquad\hat{\boldsymbol{\theta}}^{\mathrm{ref}}\in\operatorname*{argmin}_{\boldsymbol{\theta}\in\Theta_{n}(B)}\mathcal{L}_{n}(\boldsymbol{\theta};\widehat{\mathbf{P}}^{\mathrm{ref}}).$
Theorem 4 (Empirical $L^{2}$ consistency).
Suppose Assumptions 1, 2, and 7 hold. Then each optimization problem above admits a unique minimizer over $\Theta_{n}(B)$. Moreover,
$\frac{1}{n}\sum_{i=1}^{n}\big(\hat{\theta}_{i}^{\mathrm{lto}}-\theta^{\star}(X_{i})\big)^{2}=O_{\mathbb{P}}\!\left(\frac{\log n}{n}\right),$
and for an independent reference sample of size $m$,
$\frac{1}{n}\sum_{i=1}^{n}\big(\hat{\theta}_{i}^{\mathrm{ref}}-\theta^{\star}(X_{i})\big)^{2}=O_{\mathbb{P}}\!\left(\frac{1}{n}+\frac{\log n}{m}\right).$
Here $A_{n}=O_{\mathbb{P}}(a_{n})$ means that $A_{n}/a_{n}$ is bounded in probability.
Theorem 4 establishes that the empirical projection consistently recovers the population calibrated score evaluated at the observed sample points. The resulting error can be decomposed into two components: the statistical error from estimating the pairwise preference matrix, and the intrinsic $n^{-1}$ term arising from the strong convexity of the empirical logistic objective. More generally, for any empirical preference matrix $\widehat{\mathbf{P}}=\{\hat{p}_{ij}\}$ satisfying $\hat{p}_{ij}\in[0,1]$ and $\hat{p}_{ij}+\hat{p}_{ji}=1$, the empirical minimizer over $\Theta_{n}(B)$ exists uniquely and satisfies
$\frac{1}{n}\sum_{i=1}^{n}(\hat{\theta}_{i}-\theta^{\star}(X_{i}))^{2}=O_{\mathbb{P}}\!\Big(n^{-1}+\|\widehat{\mathbf{P}}-\mathbf{P}\|_{\max}^{2}\Big),$
where
$\|\widehat{\mathbf{P}}-\mathbf{P}\|_{\max}=\max_{i\neq j}|\,\hat{p}_{ij}-p_{ij}\,|.$
This bound shows that the projection step itself does not introduce additional statistical complexity: the overall error is governed by the accuracy of preference estimation together with the intrinsic parametric $n^{-1}$ term. In particular, when the preference matrix is estimated sufficiently accurately, the $n^{-1}$ term dominates, yielding the optimal parametric rate for finite-dimensional M-estimation.
A further notable feature is that the rate is dimension-free, depending only on sample size and preference-estimation error rather than on the ambient dimension or geometric complexity of the observation space. This highlights a key advantage of the proposed framework in high-dimensional or non-Euclidean settings, where many geometric or depth-based procedures suffer from dimensional degradation.
Remark 5 (Spectral approximation).
Section 2.2 also introduced a spectral estimator based on the stationary distribution of a Markov operator constructed from $\widehat{\mathbf{P}}$. This estimator can be viewed as a linearized approximation to the logistic M-projection studied above. A separate finite-sample analysis of the spectral estimator is beyond the scope of the present paper.